### v1

# Cross-Embodied Co-Design for Dexterous Hands

Abstract
Dexterous manipulation is limited by both control and design, without consensus as to what makes manipulators best for performing dexterous tasks. This raises a fundamental challenge: how should we design and control robot manipulators that are optimized for dexterity? We present a co-design framework that learns task-specific hand morphology and complementary dexterous control policies. The framework supports 1) an expansive morphology search space including joint, finger, and palm generation, 2) scalable evaluation across the wide design space via morphology-conditioned cross-embodied control, and 3) real-world fabrication with accessible components. We evaluate the approach across multiple dexterous tasks, including in-hand rotation with simulation and real deployment. Our framework enables an end-to-end pipeline that can design, train, fabricate, and deploy a new robotic hand in under 24 hours. The full framework will be open-sourced and available on our website.
Figure 1: We present our cross-embodied co-design framework for dexterous hands that jointly optimize control policies and hardware design. Four co-design hands are deployed sim-to-real.
1 Introduction
Endowing robots with human-level dexterity critically requires both the development of fine control (Wang et al., 2025; Qi et al., 2023; Zeng et al., 2019) and dexterous hardware design (Zorin et al., 2025; Shaw et al., 2023; Unitree Robotics, ; Allegro Hand, ; Christoph et al., 2025). Traditionally, robot manipulation has been developed by decoupling mechanical design from control, with reinforcement learning and control frameworks developed on fixed hardware architectures. This separation can limit dexterity - as optimal control policies may be fundamentally constrained by the hand’s morphology, degrees of freedom, and sensing capabilities (Yin et al., 2025; Calandra et al., 2018). Co-design aims to overcome this limitation by simultaneously adapting design and control (Cheney et al., 2013; Pathak et al., 2019; Bai et al., 2025).
Co-design for manipulation especially suffers from a two fold curse of dimensionality, with a massive search space in both complex hardware design and fine dexterous control. As a result, jointly creating both requires efficiently iterating on hardware designs and quickly developing fine control. To achieve speed without loss of design complexity, prior works have often leveraged fast sampling based controllers (Williams et al., 2016; Zhao et al., 2020a; Xu et al., 2021c). However, there are challenges in using these controllers for manipulation due to sparse rewards, high contact solution space, and difficulty finding solutions farther away from initial conditions (Appendix Fig. 1) Rozzi et al. (2024); Cheng et al. (2021). As a result, co-design methods for manipulation have a limited search space (Mannam et al., 2023; Yi et al., 2025a; Schaff et al., 2022) or avoid sparse, long horizon rewards altogether by addressing locomotion only (Dinev et al., 2022; Yuan et al., 2021; Zhao et al., 2020a). We solve the scalability issue of complex design and fine control for dexterity by leveraging pre-training with a morphology-conditioned cross-embodied policy applied to generated robot hands.
Additionally, previous works (Yuan et al., 2021; Lu et al., 2025; Carlone & Pinciroli, 2019; Wang et al., 2023a; Cheney et al., 2013) have often been deployed only in simulation. The sim-to-real gap remains for many works due to over simplifications in hardware manufacturing constraints Aljalbout et al. (2025), need for generalization of simulated control to real world dynamics Huang et al. (2023), and modeling of physical properties Zhao et al. (2020b). To overcome this, we create a modular hardware platform with realistic components in simulation, enabling a feasible search with designs for sim-to-real.
Cross-embodiment co-design framework: We enable quick and efficient evaluation of co-designed robot hands through incorporating a cross-embodiment control policy into our co-design framework, allowing design to sim-to-real in under 24 hours. Modular Robot Hand We create a robot hand platform with modular mechanical and electrical structure. This allows exploration of variable kinematics, degrees of freedom, and morphology. We show this by deploying four different robot hand embodiments sim-to-real. Accurate Robot Hand Generation We create modular and physically realistic grammars and rule structures that generate robot hands that can be built and deployed sim-to-real. Design Analysis: We expand co-design beyond primitive parameters for manipulation and identify design parameters of impact for future dexterity research.
The result is a cross-embodied co-design framework with efficient evaluation and generation of robot hands. We demonstrate the effectiveness of this approach through sim-to-real with in-hand rotation across multiple generated designs. More result videos can be found at: https://an-axolotl.github.io/co-design-for-dexterity.github.io/
2 Problem Setting and Notations
We aim to jointly optimize robot embodiments and control, capturing their fundamental interdependence and enabling the discovery of designs where structure and control are naturally aligned. We denote the design space as $\mathcal{G}$ and the control policy as $\Pi$. Specifically, our goal is to co-design both the robotic hand morphology and the policy $(G\in\mathcal{G},\pi\in\Pi)$ such that the reward of a given manipulation task is maximized.
Morphology Representation. We formulate a robot hand design as a graph $G=(V,E,X_{v},X_{e})$. The nodes $v\in V$ represent finger segment links. The edges $e\in E$ represent kinematic joints between finger components. The maximum number of fingers used in this work is five, though the method could be expanded beyond that. We denote the number of fingers to be the width of the root node, i.e. $f=\mathrm{width}(v_{root})=\{1,\cdots,5\}$, and the length of each finger to be $l=\mathrm{depth}(v_{root,i})=\{1,\cdots,4\}$. Each finger may be of a different length, depending on the depth of its respective branch.
Graph Neural Networks. To consider the isomorphism of different graphs, we encode each $G$ with a message-passing GNN $f_{\phi}$. Given $(V,E,X_{v},X_{e})$, $f_{\phi}$ performs $K$ rounds of message passing to produce a fixed-dimensional embedding $y(G)=f_{\phi}(G)\in\mathbb{R}^{d}.$ This output $y(G)$ is then used as a design condition for subsequent control policies and evaluations.
Design-Conditioned Control Policy. We model control of our dexterous manipulation setting as a Markov Decision Process (MDP) $\mathcal{M}_{G}=(\mathcal{S},\mathcal{A},T,R,\gamma),$ where the dynamics $T_{G}$ depend on the chosen morphology $G$. The state space $s\in\mathcal{S}$ is defined as $s=\big(q,\,p_{o},\,y(G)\big),$ where $q\in\mathbb{R}^{|E|}$ are the robot joint states, $p_{o}\in\mathbb{R}^{6}$ is the object pose, and $y(G)\in\mathbb{R}^{d}$ is the graph encoding of the design $G$, where $d=F_{\max}\times(1+2L_{\max})$, corresponding to up to $F_{\max}=5$ fingers, each with at most $L_{\max}=3$ links.
The action space $\mathcal{A}$ consists of actuator commands of the joint positions. Since not all actuators are valid for all morphologies, we incorporate an action mask $m(G)\in\{0,1\}^{|\mathcal{A}|}$ and apply it to the raw policy output:
$a_{t}=\pi_{\theta}(s_{t})\odot m(G),$
where $\odot$ denotes element-wise multiplication. This ensures that the design-conditioned policy $\pi_{\theta}:s_{t}\rightarrow a_{t}\in\Pi$ can be shared across designs without producing invalid actions. The transition function $T(s_{t+1}\mid s_{t},a_{t})$ represents the physics integration in simulation. The reward $R(s_{t},a_{t})$ is task-specific and includes object states, joint efforts, object proximity to the hand, and more, depending on the tasks. The objective of policy learning for a fixed $G$ is to maximize expected discounted return:
$J(\pi,G)\;=\;\mathbb{E}_{\tau\sim\pi,\mathcal{M}_{G}}\Big[\sum_{t=0}^{K}\gamma^{t}R(s_{t},a_{t})\Big]$ (1)
where $\tau=(s_{0},a_{0},s_{1},\dots)$ denotes a trajectory induced by $\pi$ in $\mathcal{M}_{G}$.
Co-Design as Bi-Level Optimization. The co-design problem seeks to jointly optimize morphology and control. Formally, the optimal design $G^{*}$ and its corresponding policy $\pi^{*}$ are defined by
$(G^{*},\pi^{*})=\arg\max_{G\in\mathcal{G},\,\pi\in\Pi}\;J(\pi,G)$ (2)
Equivalently, this can be written as a bi-level optimization problem:
$G^{*}$ $=\arg\max_{G\in\mathcal{G}}\;J(\pi_{G}^{*},G),$ (3)
$\pi_{G}^{*}$ $=\arg\max_{\pi\in\Pi}\;J(\pi,G).$ (4)
Here, the inner loop learns the best control policy for a given morphology, while the outer loop searches for the morphology whose optimal policy yields the highest task performance. This joint design-control space can be high-dimensional and computationally intractable. We introduce our method of first training a cross-embodiment base policy $\tilde{\pi}_{\theta}$ from sampled designs $G\sim p(G)$, and then fine-tuning on generated designs to help explore both the design and control space efficiently.
3 Method
Our approach enables sim-to-real robot hand co-design through cross-embodiment learning across embodiment families. An overview of our method is shown in Figure 2. We first generate a set of different embodiments by randomly sampling different design rules. Rather than learning separate policies for each hand design, we group morphologies into families based on shared kinematic structures, asymmetric or symmetric, and variable finger counts—enabling efficient generalization across fingertip types, degrees of freedom, link lengths, palm geometries, and finger placements within each family. To make the control policy aware of the actual morphology, while considering the isomorphism of design, we condition the control policy on encoded morphology. Our physical grammar system reflects real hardware constraints through pre-computed collision geometries, joint limits, and manufacturable component specifications, ensuring designs transfer effectively to physical systems.
Components. The framework consists of four main components:
Generation $G\sim p_{grammar}(r,c)\quad\triangleright\text{Procedural constraint hand generation}$ (5)
Design encoder $g=h(G)\quad\triangleright\text{Morphology embedding for policy conditioning}$ (6)
Cross-embodiment policy $\pi_{\theta}(a|s,g)\quad\triangleright\text{Morphology-conditioned control}$ (7)
Design value network $V_{design}(g)\quad\triangleright\text{Learned design evaluation for search guidance}$ (8)
where $g$ is the graph embedding of morphology $G$, $s$ and $a$ are states and actions, and the grammar generator creates diverse hand designs through procedural rules encoding palm geometry, finger count $f\in\{3,4,5\}$, joint configurations $j_{f}\in\{0,2,3\}$ per finger, link lengths (1-10 per joint), and fingertip types. These grammars are pre-designed real sub-components with realistic collision geometries, manufacturing constraints, and modular connections (Appendix Fig. 2). During co-design, the system iteratively (i) generates candidate morphologies through Graph Heuristic Search (Zhao et al., 2020a), (ii) evaluates designs using the pre-trained cross-embodiment policy, and (iii) manufactures designs for real testing.
Figure 2: Overview of our method. Stage 1: We first randomly sample embodiments, including morphology and degrees of freedom. We then pre-train a morphology-conditioned policy across embodiments. Stage 2: The design is generated using modular grammars, with assembly rules selected based on prior performance. The cross-embodied policy is then deployed to evaluate different designs in simulation. Stage 3: Best designs and fine tuned control are then manufactured and deployed on real hardware across unseen objects.
3.1 Model Objective
Design evaluation leverages cross-embodiment to assess morphology performance. The design value network learns from observed design-performance pairs:
$\mathcal{L}_{design}=\mathbb{E}_{d\sim\mathcal{D}_{evaluated}}\left[(V_{design}(G_{d})-\mathcal{T}(d))^{2}\right]$ (9)
where $\mathcal{T}(d)$ represents the maximum observed task performance for design $d$ using the cross-embodiment policy. The search explores the design space through Graph Heuristic Search (Zhao et al., 2020a) with $\epsilon$-greedy expansion, maintaining a lookup table $\mathcal{T}:\mathcal{D}\rightarrow\mathbb{R}$ that maps designs to rewards and guides value network training. At each iteration, $K$ candidate designs are generated either through random grammar rule application (exploration) or value-guided expansion from promising partial designs (exploitation), with the lookup table updated via $\mathcal{T}(d)=\max(\mathcal{T}(d),R(G))$ for newly evaluated designs.
3.2 Architecture and Training
Hand Generation for Cross-Embodiment Pre-training. We develop a parametric hand generator that constructs diverse, physically-valid morphologies through a two-stage process. First, the generator samples palm layouts by placing finger bases on a variable circular radius under configurable placement modes (symmetric, asymmetric, or anthropomorphic), discretizing the circle into slots with minimum angular separation enforced via rejection sampling. A convex palm mesh is constructed by computing a 2D convex hull of grip points around each motor position and extruding it to a specified thickness. Second, each finger’s internal morphology is independently sampled from a structured design space controlling: (1) the number of actuated joints (2–3 servos beyond the base), (2) two discrete “grammar codes” $g_{1},g_{2}\in\{1,\ldots,10\}$ that determine proximal and distal segment lengths through link scaling, and (3) fingertip geometry (standard, wedged, rounded, or thin variants). This produces thousands of hand variants by sampling from uniform distributions across grammar parameters with fully randomized designs. Each generated hand includes pre-computed collision geometries via CoACD convex decomposition for all modular components, realistic joint limits derived from CAD specifications, and consistent kinematic representations.
Cross-Embodiment Policy Pre-training. We group generated hands into kinematic families based on structural similarity—symmetric radial configurations, anthropomorphic arrangements, and variable finger counts—enabling targeted policy learning within each family. The cross-embodiment policy uses PPO enhanced with morphology-conditioning:
$\mathcal{L}_{policy}=-\mathbb{E}_{(s,g)}\left[\min\left(\frac{\pi_{\theta}(a|s,g)}{\pi_{old}(a|s,g)}\hat{A}_{t},\text{clip}\left(\frac{\pi_{\theta}(a|s,g)}{\pi_{old}(a|s,g)},1-\epsilon,1+\epsilon\right)\hat{A}_{t}\right)\right]$ (10)
where $g$ represents the morphology embedding produced by our graph encoder. Each family trains separately on 2000-8000 hand variants, with the policy learning unified control strategies that generalize across finger counts, joint configurations, and geometric variations within the family. The morphology embedding $g$ enables the policy to adapt its control strategy based on the specific kinematic structure and available degrees of freedom.
Graph Heuristic Search Algorithm. Following cross-embodiment pre-training, we employ Graph Heuristic Search to optimize hand designs for specific tasks. The algorithm iteratively generates candidate designs through $\epsilon$-greedy expansion from a design value network $V_{design}(g)$ that predicts task performance without full evaluation. At each iteration, we generate $K$ candidate morphologies, evaluate them in parallel using the pre-trained cross-embodiment policy, and update both the lookup table $\mathcal{T}:\mathcal{D}\rightarrow\mathbb{R}$ and value network based on observed performance. The search maintains design diversity through grammar rule application while focusing computational effort on promising regions of the morphological space.
Design Optimization The search begins from an initial design with a randomly configured palm layout and randomized finger bases, marking fingers as non-terminal nodes requiring expansion. At each iteration, the algorithm generates candidates through sequential finger-by-finger expansion: for each non-terminal finger, it evaluates all valid parameter choices by constructing successor designs and scoring them with the GNN, which predicts performance from graph representations augmented with Gumbel noise ($\text{noise\_scale}\approx 0.4$) for exploration. The highest-scoring successor is selected, repeating until all fingers are finalized into complete designs. These candidates are evaluated in parallel simulation across multiple hand groups, yielding task-specific performance scores. The lookup table is updated with each complete design’s score and its partial ancestors’ scores (enabling credit assignment during construction), using effective keys that account for morphological equivalences (finger permutations in symmetric layouts, thumb slot variations in anthropomorphic layouts). The GNN is trained via supervised regression on this growing dataset, minimizing squared error between predicted and observed scores with L2 regularization and gradient clipping. This updated value network guides subsequent candidate generation with refined predictions, while an annealing epsilon schedule ($0.4\rightarrow 0.05$) shifts from exploration to exploitation, and path-level tabu tracking prevents revisiting identical construction sequences within each batch, ensuring sustained diversity.
Sim-to-Real Transfer. The best-ranked design from search is fine-tuned with domain randomization over actuator characteristics, contact and friction parameters, and object poses to improve robustness (Andrychowicz et al., 2020; Tobin et al., 2017). For deployment, we train the controller as a blind policy, i.e., removing object-state inputs, to match onboard sensing and reduce sim–to-real mismatch. There is no camera or tactile feedback in the system. The cross embodied policy for the respective design family is fine tuned by loading the checkpoint and tuning weights for the stateless observations. Manufacturing the real robot hand follows the same programmatic grammar used in simulation: the best selected design graph is directly converted to modular hardware specifications with 3D-printed components and precisely matched actuator constraints. A mounting interface is added to the palm, each joint and links are directly 3D printed, and then assembled with Dynamixel actuators. After tuning PID on hardware, the blind policy is deployed on the physical system.
4 Experiments
Figure 3: Left to Right. 1) Variable finger length of 1-10 using flat modular stacks. 2) Example embodiment with varying degrees of freedom with fingers of 3 or 4 servo motors. 3) Fingertip Variations 4) Finger number and example embodiment variations.
We evaluate our co-design framework on three dexterous manipulation tasks, including one with simulation and real-world deployment for in-hand rotation using generated designs. Our experiments demonstrate that cross-embodiment learning paired with modular hardware graph grammars enables efficient and generalizable co-design for manipulation without sacrificing the runtime or dexterity commonly lost when using traditional co-design control methods. We show successful performance across multiple tasks and demonstrate effective sim-to-real transfer.
Environments. We test the co-design framework across three manipulation tasks, including sim-to-real in-hand rotation. These tasks comprise in-hand rotation, grasping on a table top, and object flipping about the z axis on a table top with fifteen randomized objects. We run 50 iterations with 40 designs at each iterations. Each design is tested across 128 randomized environments including randomized joint positions, physical parameters, and object type. Please see 5 for setup.
Table 1: Comparison of best designs continuous in-hand rotation in rad/s. Baselines search across 2000 designs for each with stronger reward for manipulation (Appendix Baselines). For the LEAP hand we use the Leap-Reorientation environment and train their default policy code and object (1 object). Then we retrain the policy on our randomized object dataset, blind which fails to rotate.
Method Run Time (hours) Cont. Ang. Velocity (rad/s)
Ours 6.48 3.3
Ours w/o fine tuning 5.18 1.85
Ours w/ MPPI 20.0 0.62
Leap, Single Cube w/ Vision 2.0 0.47
RoboGrammar 23.0 0.26
Monte Carlo 15.2 0.20
Blind Leap Hand 2.0 0.0
Baselines. We compare our co-design framework against the following baselines for in-hand rotation: (1) RoboGrammar (Zhao et al., 2020a), a procedural generation algorithm using graph-based design rules and local search optimization for robot morphology design; (2) Monte Carlo Tree Search, a tree-based exploration method that systematically searches the design space by building a search tree and using random rollouts for evaluation; (3)Our method with MPPI control (4)the LEAP hand with vision trained on a single object trained with PPO Shaw et al. (2023)(5) A blind LEAP robot hand with our randomized objects for comparison to standard hardware.
Compared to these baselines, our method achieves finer control and is able to successfully complete manipulation tasks. Robogrammar achieves a best design rotation speed of 0.26 rad/s and a time to fall of 4.63 seconds for the in-hand rotation task. By contrast, our best designs without fine tuning achieves a 1.85 rad/sec in simulation with no time to fall over the evaluation time of three minutes. Monte Carlo performs worse than the graph heuristic search, with a time to fall of 2.71 seconds. We find reinforcement learning is substantially better suited for manipulation co-design tasks as it can better overcome sparse rewards, long horizon tasks, and achieve complex finger gaits needed for true dexterity.
Table 2: End-to-end pipeline timing for generating, manufacturing, and deploying one robot hand sim-to-real.
Component Time (hours)
3D Printing 12.0
Algorithm 6.48
Assemble 0.8
Sim-to-Real 2.0
Sim-to-Real Transfer. To demonstrate our framework’s real-world applicability, we conduct zero-shot blind deployment from simulation to reality. We use the best found design to deploy a blind policy with our random rotation objects within our domain randomization constraints. Three co-design embodiments are shown to show hardware tractability of our modular setup and show design tradeoffs. One baseline anthropomorphic hand is also deployed sim-to-real for comparison. We deploy the policy zero shot, open loop with no state feedback of the object type and no object position. Only encoder information of servo motors and the morphology encoding are used as the observation.
5 Results: Analysis of Generated Robot Hands
Figure 4: Top Left: Three finger co-design hand, best design found by algorithm. Top Right: Five finger symmetric co-design hand. Bottom Left: Four finger co-design hand with thin fingertips. Bottom Right: Five finger anthropomorphic baseline.
5.1 Real World Deployment
To show our method is able to be built and transferred to the real world, we tested seventeen unseen objects with variable friction, compliance, and shape. A blind policy was trained in simulation on a subset of high performing designs found by rotation simulated environment. We compare three co-design hands, including the best found optimal 3-fingered hand, to an anthropomorphic baseline. We find robot hands with radial symmetry greatly outperform anthropomorphic robot hands on blind in-hand rotation. We choose these designs to show the impact of morphology on task success and the ability to accurately predict task success ranking in real. From simulation results, the anthropomorphic robot hand trained across the randomized objects achieves a 0.73 rad/s in simulation. The three fingered robot hand achieves 1.85 rad/sec without fine tuning. The robot hands were deployed with the morphology-conditioned cross-embodiment policy. The inputs are the morphology’s encoding and the joint state positions only, with no state information, while the robot hands ran across objects until either reaching a minute of continuous rotation or failure.
Table 3: Four designs are shown in order of simulation success ranking showing accuracy of our algorithm evaluation sim-to-real. Task performance (seconds) across hand types and objects. A dash indicates not achieved.
Hand type Pink polygon Tennis ball Rubik’s cube Green block Purple ball Star fruit Pink disk Peach can Yel. block
3-finger, optimum 9 13 9 9 5 15 10 6
5 Finger 12 18 8 14 18
4 Finger 10 9
Anthropomorphic 25 13
We show the time used to rotate a given object 360 degrees, and a subset of the results is shown in Table 3. Failure to rotate within one minute is denoted as a dash. The three fingered robot hand executed grasps which were able to generalize more effectively to unseen objects. The anthropomorphic and four finger robot hands were able to blindly rotate a small subset of objects - three out of seventeen tested objects, as predicted to be lower performing by our algorithm. Objects often became stuck between finger gaits. Across the same seventeen objects, the three fingered robot hand generalized substantially better generalization. Across all seventeen objects, only two failed with one two small between the robots bling gait to reach. The three fingered morphology generalized across irregular objects including pine cones, pentagons, and soft objects. The gait of the three fingered robot hand is three altering diagonals, allowing the object to be rotated with little displacement. Please see videos for details.
5.2 Simulation Task : Design Results
In simulation, we deploy across three tasks including grasping, in hand object rotation, and object flipping. Grasping Grasping is lifting the randomized object to the palm with a stable grasp and fixed wrist. Across randomized objects, thinner fingertips were primarily picked with 54% of designs having this feature. All best designs had full degrees of freedom per finger (4 servo motors) and five fingers of symmetric and anthropomorphic layouts. These designs complete a successful grasp without dropping in under a half second. Rotation In hand rotation had the highest standard deviation, with 25th percentile hands having less than 25 degrees of rotation per second and the top 75th percentile hands more than double (67 deg/sec). Best designs used three fingers achieving - without fine tuning - 1.85 rad/sec across objects and 3.3 rad/sec with fine tuning. Standard fingertips worked best with 64% of designs using this attribute. Thinner fingertips resulted in unstable rotation. Flipping Flipping is composed of rotating the object around the z axis underhand and requires forward rotation with passing the object back to a start position for additional rotation. Best designs (58% of all generated designs) used wedged or thin fingertips on one side to lift the object easily and standard fingertips on the otherside of the palm to have controlled reset with four and five finger symmetric hands. Across all tasks, full degrees of freedom (4 servos per finger) consistently performed asymmetric designs, including those with thumb layouts. As the hardest task due to a fixed wrist and variable object heights, best designs achieve flipping in less than 6 seconds.
Figure 5: Top Left: Grasping task with hold time over a minute and quick grasps (<< 0.1 sec).Bottom: Best found design of flipping task where an object rotates along the z-axis using a fixed wrist. Bottom Right: Normalized reward of best design for Flipping, Grasping, and In-Hand Rotation.
6 Why Morphology?
We conduct a parameter analysis for robotic hand manipulation by systematically sampling across 12 critical hardware categories in the Isaac-Reorient-Cube-Leap task on a LEAP hand (Shaw et al., 2023). Our sampling approach used uniform grid sampling with 100 samples per parameter across their physically meaningful ranges (Appendix Fig. 14-16), covering three parameter categories: physics properties (static/dynamic friction, restitution), actuator characteristics (stiffness, damping, effort limits, velocity limits), and morphological scaling factors (palm width, finger length, mass scaling). Each parameter class was trained using PPO, yielding a dataset of 1,200 training runs.
Figure 6: Correlation of physical parameters to improving rotation from Bayesian Sampling on a LEAP hand. Parameters of highest impact are morphology and control parameters.
Our analysis (Fig. 5 and Appendix Grid Sampling) reveals four parameters with particularly strong effects on manipulation performance: finger body length scale showed the strongest positive correlation (r=0.748), while palm width scale exhibited a strong negative correlation (r=-0.729), suggesting that wider palms impede dexterous rotation. Of these, morphology parameters had the highest positive impact. Additionally, two motor control characteristics, damping (r=-0.476) and dynamic friction (r=-0.459), showed moderate negative correlations with performance. This provides a strong argument for co-design since the top four parameters can be optimized through control and morpholgoy. Material properties were found to be of lower impact. A non-linear relationship is observed through polynomial curve fitting indicate that optimal manipulation performance occurs within specific parameter ranges rather than at extremes. Morphology and conrol having the most pronounced impact on task success compared to material or contact properties.
7 Related Works
Co-Design of morphology and control policy. Traditional design approaches first fix hardware architecture, then optimize control, co-design treats morphology and policy as a joint space that often outperforms separate optimization (Ha et al., 2016; Schaff et al., 2019; Matthews et al., 2023). However, the morphology-policy joint optimization is high-dimensional (Chen et al., 2021). Previous works focused on optimizing the design only (Ha et al., 2021; Xu et al., 2024; Kodnongbua et al., 2023) based on a fixed control policy, or co-design with a limited design space (Luck et al., 2020; Islam et al., 2024; Yi et al., 2025b; Dong et al., 2024). Most co-design work has remained in simulation due to challenges in sim-to-real transfer of both control policy and manufacturable designs (Xu et al., 2021a; Xiong et al., 2023; Dong et al., 2023a). While most works focused on locomotion (Yuan et al., 2021; Wang et al., 2023c; Dong et al., 2023b), manipulation presents unique challenges due to the large hardware design space (Piazza et al., 2019), fine control for contact-rich dynamics (Qi et al., 2023; Yuan et al., 2020; Zakka et al., 2023). Evolutionary algorithms scale poorly to complex morphologies(Lipson & Pollack, 2000; Cheney et al., 2013). Programmatic spaces, as in Zhao et al. (2020a), add compositional rules for valid topologies, improving sample efficiency and structured exploration. We combine programmatic design with cross-embodiment learning to efficiently search the joint space without sacrificing fine motor control.
Cross-Embodiment Learning Learning policies that generalize across different robot morphologies addresses the sample efficiency limitations of training separate models for each embodiment (Doshi et al., 2024; Gupta et al., 2021; O’Neill et al., 2024). Current approaches leverage shared neural network architectures with morphology-conditioned inputs (Huang et al., 2020; Attarian et al., 2023; Furuta et al., 2022), function approximators that encode embodiment parameters (Zhao et al., 2020a), morphological pretraining (Li et al., 2024; Strgar & Kriegman, 2025), and meta-learning frameworks for few-shot adaptation to new morphologies (Nagabandi et al., 2018; Wang et al., 2023b). Graph neural networks have shown particular effectiveness for variable-topology robots by learning permutation-invariant representations over robot joints and links, enabling zero-shot transfer to previously unseen kinematic structures (Wang et al., 2018; Liu et al., 2023; Lu et al., 2025). However, transfer performance degrades significantly with large morphological differences, particularly in contact-rich manipulation tasks where small design variations can drastically affect policy performance (Doshi et al., 2024).
Sim-to-Real Transfer Distribution shift between simulated training and real-world deployment remains challenging - especially for co-designed systems (Chen et al., 2021; Schaff et al., 2023). The reality gap stems from complex physical dynamics, control noise, and manufacturing tolerances that compound when both morphology and control are optimized in simulation (Peng et al., 2018; Muratore et al., 2022) Current solutions include system identification techniques that adapt learned policies to real hardware through online fine-tuning (Du et al., 2021). While recent work combines reinforcement learning with differentiable optimization for locomotion (Xu et al., 2021b), co-design for sim-to-real and dexterous manipulation remains largely unexplored.
8 Lessons Learned and Limitations
Through our experiments on cross-embodied co-design for three dexterous manipulation tasks, we find the following: 1) Cross-embodiment enables tractable co-design scaling. Cross-embodiment evaluation provides a practical approach to assess designs for complex manipulation without sacrificing fine control for computational speed, making co-design optimization feasible for real-world deployment. 2) Realistic grammar-based design enables zero-shot transfer. Grounding design grammars in physical components allows for seamless sim-to-real transfer without additional fine-tuning, as the modular structure maintains consistent physical properties across domains. 3) Task-specific optimization can outperform anthropomorphic design. Our results demonstrate that morphologies optimized for specific manipulation tasks for flipping and rotation tasks. Further research in general non-anthropomorphic hands can be explored. 4) Morphology dominates the design space. Among all design parameters, morphological choices have the largest impact on manipulation performance, suggesting that shape and structure optimization should be prioritized over other design considerations.
Limitations. While our framework demonstrates effectiveness across multiple manipulation tasks, several key bottlenecks remain. Our current approach only recomputes palm geometry while keeping other components modular and pre-defined, limiting full exploration of the morphological design space. Additionally, individual designs remain optimized for specific tasks rather than generalizing across multiple manipulation skills—incorporating multi-task design averaging or weighting mechanisms represents a critical direction for future work. We also lack systematic analysis of design diversity and convergence properties in our optimization process, which could lead to mode collapse or insufficient exploration. Finally, our framework focuses primarily on morphological optimization while neglecting other critical parameters such as compliance, material properties, and actuation schemes that could further improve manipulation performance.

### v2

@@
 Our framework enables an end-to-end pipeline that can design, train, fabricate, and deploy a new robotic hand in under 24 hours.
-The full framework will be open-sourced and available on our website.
+Results and build guide are available at: https://an-axolotl.github.io/co-design-for-dexterity.github.io/
 Figure 1: We present our cross-embodied co-design framework for dexterous hands that jointly optimize control policies and hardware design. Four co-design hands are deployed sim-to-real.
@@
 The cross-embodied policy is then deployed to evaluate different designs in simulation.
-Stage 3: Best designs and fine tuned control are then manufactured and deployed on real hardware across unseen objects.
+Stage 3: Best designs and control pairs are then manufactured and deployed on real hardware across unseen objects.
 3.1 Model Objective
@@
 Sim-to-Real Transfer.
-The best-ranked design from search is fine-tuned with domain randomization over actuator characteristics, contact and friction parameters, and object poses to improve robustness (Andrychowicz et al., 2020; Tobin et al., 2017).
-For deployment, we train the controller as a blind policy, i.e., removing object-state inputs, to match onboard sensing and reduce sim–to-real mismatch.
+The best-ranked design from search is selected and tuned alongside domain randomization over actuator characteristics, contact and friction parameters, and object poses to improve robustness (Andrychowicz et al., 2020; Tobin et al., 2017).
+We deploy the design conditioned controller as a blind policy, i.e., removing object-state inputs, to match onboard sensing and reduce sim–to-real mismatch.
 There is no camera or tactile feedback in the system.
-The cross embodied policy for the respective design family is fine tuned by loading the checkpoint and tuning weights for the stateless observations.
 Manufacturing the real robot hand follows the same programmatic grammar used in simulation: the best selected design graph is directly converted to modular hardware specifications with 3D-printed components and precisely matched actuator constraints.
@@
 The gait of the three fingered robot hand is three altering diagonals, allowing the object to be rotated with little displacement.
-Please see videos for details.
+Cross embodied policies also lead to improved generalization and dexterity over PPO trained for a single hand (Appendix 9.1.2).
+Please see video and website for details.
 5.2 Simulation Task : Design Results